OpenStudy (anonymous):

Which integral expression can be used to find the area in the first quadrant bounded below by the x-axis and above by the curves y = sin x and y = cos x?

OpenStudy (anonymous):

In the first quadrant, the graphs of sin(x) and cos(x) intersect at x = pi/4. In that interval cos(x) has larger values than sin(x) so the integral will be for cos(x) - sin(x) and its limits will be x = 0 to x = pi/4 \[\int\limits_{0}^{\pi/4}\cos(x)-\sin(x)dx\]

OpenStudy (anonymous):

I am not sure which one it is.

OpenStudy (anonymous):

a = 0, b = pi/2... Upper limit is cos x and lower limit is sin x... Therefore... \[\int\limits_{0}^{\pi/2} (\sin x - \cos x )dx\]

OpenStudy (anonymous):

Oh okay I misinterpreted it sorry! The question is asking you about area above x-axis in the first quadrant bounded by sin and cos graphs. The graph looks like this